O You are looking at an investment that will pay $1200 in 5 years if you invest $1000 today. What is the rate of interest?
Answer #1
Rate of interest = 3.71%
Dear Student
Thank you for using Chegg
Please find below the answer
Statementshowing Computations
Paticulars
A = P(1+r/100)^n
1200 = 1000(1+r/100)^5
1.2 = (1+r/100)^5
1.2^(1/5) = 1+r/100
1.0371 = 1+r/100
.0371 = r/100
r = 3.71%